OpenStudy (anonymous):

I need help with six questions just like the following. Can u explain HOW to do it? Pls HELP! MEDAL!!!! 2(x+12)-2x=3(x-2) A. One real solution B. Infinite real solution C. No real solution

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

not really

OpenStudy (texaschic101):

All it means is take the number outside of the parenthesis, and multiply it by everything in the parenthesis...let me show you.. 2(x + 12) is the same as 2 * x + 2 * 12 = 2x + 24 do you understand that ?

OpenStudy (anonymous):

Oh yeah, that's easy! Okay, what next?

OpenStudy (texaschic101):

2(x + 12) - 2x = 3(x - 2) --- distribute through the parenthesis 2x + 24 - 2x = 3x - 6 are you still following me ? Any questions so far ?

OpenStudy (anonymous):

Yup, still following you. No questions.

OpenStudy (texaschic101):

2x + 24 - 2x = 3x - 6 --- combine any like terms that you have 24 = 3x - 6 -- you see how the 2x's cancelled now we want to get x by itself, so we will add 6 to both sides 24 + 6 = 3x - 6 + 6 --- the 6's on the right side cancel 30 = 3x -- now divide both sides by 3 30/3 = x 10 = x still following ? Ask questions if you have any ?

OpenStudy (anonymous):

distribute then combine like terms then simplify

OpenStudy (texaschic101):

basically...yes. And since we found x, there is only 1 solution

OpenStudy (texaschic101):

IF your answer would have been : 3x - 2 = 3x - 2...then that would be infinite solutions...and if it would have come out like 3 = 2...that would be no solutions

OpenStudy (anonymous):

YES! That's what i guessed. Oh wait, I forgot to post that reply. :( Oops, I meant to tell you that one solution was my guess. I see what you did but now I'm gonna try another one on my own and then you can check it…..pls?

OpenStudy (texaschic101):

I will check it for you ?

OpenStudy (anonymous):

we are here to help! :)

OpenStudy (anonymous):

Thanks @icecreamman and @texaschic101. Gimme a couple secs to work it out

OpenStudy (texaschic101):

icecreamman...cool name

OpenStudy (anonymous):

Ikr!

OpenStudy (anonymous):

Okay, so I think the equation 5(x+8)-7x=-(2x+4) is a no solution

OpenStudy (texaschic101):

5(x + 8) - 7x = -(2x + 4) --- distribute 5x + 40 - 7x = -2x - 4 -- combine like terms -2x + 40 = -2x - 4 you are correct...there is no solution because you cannot put the same number for x in this problem and get a correct answer...good job

OpenStudy (anonymous):

YAY! Okay, sorry but I have a couple more I need you to check. I think the equation 24x+12=2(12x+6) is a infinite real solutions

OpenStudy (texaschic101):

24x + 12 = 2(12x + 6) -- distribute 24x + 12 = 24x + 12 yep..right again...infinite solutions

OpenStudy (anonymous):

I'm on a roll. Lol. Okay, this is about to be rapid fire. Ready?

OpenStudy (texaschic101):

I am sooo ready...go

OpenStudy (anonymous):

icecreamman would like to thank you @texaschic101

OpenStudy (texaschic101):

for what ?

OpenStudy (anonymous):

I think 8(x+7)=6(x-4) is a one real solution

OpenStudy (anonymous):

for saying that icecreamman has a cool name :) @texaschic101

OpenStudy (anonymous):

And so is 5x-20=3(x-4). Thats a one real solution too

OpenStudy (texaschic101):

8(x + 7) = 6(x - 4) 8x + 56 = 6x - 24 8x - 6x = -24 - 56 2x = - 80 x = -40 yep...one real solution oh, icecreamman...I just said that because it is true....I like the name...I could use some ice cream right about now...lol

OpenStudy (anonymous):

I COULD USE ICE CREAM TOOOOO! I'm supposed to be on spring break and instead I'm working on this isn

OpenStudy (texaschic101):

spring break is not for homework...its for relaxing and having fun

OpenStudy (anonymous):

Your telling me……smh. I hate my school right now.

OpenStudy (anonymous):

Did you see my other guess?

OpenStudy (texaschic101):

hold on...let me refresh my screen...seeing a lot of black diamonds

OpenStudy (anonymous):

I think it's still gonna be there. It's on mine too

OpenStudy (texaschic101):

what other guess ? I don't see it

OpenStudy (texaschic101):

oh wait.... 5x - 20 = 3(x - 4) 5x - 20 = 3x - 12 5x - 3x = -12 + 20 2x = 8 x = 4 one solution

OpenStudy (texaschic101):

anymore ?

OpenStudy (anonymous):

Last one!

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

12x+21=3(4x+6) is a no solution?

OpenStudy (texaschic101):

12x + 21 = 3(4x + 6) 12x + 21 = 12x + 18 you got it again...no solution

OpenStudy (anonymous):

YAY! Ohmygoodness you helped me soooooo much! Thanks you, thank you, thank you!!!! :)

OpenStudy (texaschic101):

lol...anytime. You seem to know what you are doing..now you will be able to help others that do not know

OpenStudy (anonymous):

Im so happy I'm finally done with that assignment. THanks for everything!